OpenStudy (anonymous):

Simplify -3(-2-3+6)-(-5(-2)+3(-2-4))

OpenStudy (anonymous):

what did you get? these are getting tougher

OpenStudy (anonymous):

btw i am wondering, are the instructions really "simplify"? it should be "compute"

OpenStudy (anonymous):

−3(−2−3+6)−(−5(−2)+3(−2−4)) −3⋅1−(5⋅2+3(−6)) Thats what i got so far. and yeah they say simplify.

OpenStudy (anonymous):

First you have *-3 to all: -3(-2-3+6) = -3 . -2 -3.-3-3.6

OpenStudy (anonymous):

k looks good so far

OpenStudy (anonymous):

i dont know what to do now?

OpenStudy (anonymous):

−3⋅1−(5⋅2+3(−6))

OpenStudy (anonymous):

\(-3\times 1=-3\)

OpenStudy (anonymous):

so now it will be -3−(5⋅2+3(−6))

OpenStudy (anonymous):

\[-3-(5\times 2+3\times (-6))\]

OpenStudy (anonymous):

then \(5\times 2=10\) and \(3\times -6=-18\)

OpenStudy (anonymous):

(+).(+)=+ (-).(-)=+ ( _ ).(+)=(-) Remember it !

OpenStudy (anonymous):

\[-3-(10-18)\]

OpenStudy (anonymous):

Ok what does that look like?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

two more steps \[10-18=-8\] \[-3-(-8)\]

OpenStudy (anonymous):

one more step \[-3-(-8)=-3+8\]

OpenStudy (anonymous):

and the final answer is 5?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

Simplify the following: -(-5 (-2)+3 (-2-4))-3 (-2-3+6) Group the negative terms in -2-4 together and factor out the minus sign. -2-4 = -(2+4): -3 (-2-3+6)-(-5 (-2)+3×-(2+4)) Evaluate 2+4. 2+4 = 6: -(-5 (-2)+3 (-6))-3 (-2-3+6) Multiply -5 and -2 together. -5 (-2) = 10: -(10-6×3)-3 (-2-3+6) Multiply 3 and -6 together. 3 (-6) = -18: -3 (-2-3+6)-(10+-18) Group the negative terms in -2-3+6 together and factor out the minus sign. -2-3+6 = 6-(2+3): -(10-18)-36-(2+3) Evaluate 2+3. 2+3 = 5: -(10-18)-3 (6-5) Subtract 5 from 6. 6-5 = 1: -(10-18)-31 Factor out a negative sign in 10-18 in order to subtract. 10-18 = -(18-10): -3×1--(18-10) Subtract 10 from 18. | 1 | 8 - | 1 | 0 | 0 | 8: -(-8)-3×1 Multiply all instances of -1 in -(-8). (-1)^2 = 1: 8-3×1 Multiply -3 and 1 together. -3×1 = -3: -3+8 Subtract 3 from 8. 8-3 = 5: Answer: | | 5

OpenStudy (anonymous):

@LadyQuietBottom:It s right ! Excelent !